An archer pulls the bowstring back for a distance of 0.39 m before releasing an arrow. The bow and string act as a spring of spring constant 400.1 N/m. At what speed will a 42.3 gram arrow leave the bow?


1
Expert's answer
2020-02-11T09:31:53-0500

From the conservation of energy:


"0.5mv^2=0.5kx^2"

"v=\sqrt{\frac{k}{m}}x"

"v=\sqrt{\frac{400.1}{0.0423}}0.39=37.93\frac{m}{s}"